OpenStudy (anonymous):

a board is cut into two pieces, one 12cm longer than another. If the uncut board is 200 cm what is the size of each cut?

OpenStudy (anonymous):

with this you can set it up as 200 = x+(x+12). this represents the uncut board of total length = the two peieces. so, 200 = 2x+12 = 188 = 2x 188/2 = 94. that is the short one, so the other one is 12 in more so it is 106.